OpenStudy (anonymous):

its 12x=-36y a direct variation?

OpenStudy (rulnick):

yes

OpenStudy (rulnick):

y = (-1/3) x

OpenStudy (anonymous):

1. -8y = 2x 3. 12x = 36y 5. y - 12 = 12x A direct variation is when an equation of a line is in slope-intercept form (y=mx+b) there is no b, or b equals 0. 1. If you simplify the equation by dividing both sides by -8 to get y alone. The equation is y= -1/4x. Since there is no b the equation is a direct variation. The constant of variation is just the number being multiplied by x. 3. To get y alone you just divide both sides by 36 and the equation in y=1/3x. There is no b so its a direct variation. X is being multiplied by 1/3 so the constant of variation is 1/3. 5. If you solve the equation for y so its in slope-intercept form its y= 12x+12. Since there is a 'b' the equation is not a direct variation. Source(s):